Predict the molecular shape of methane, the carbon
Answer #1
Concepts and reason

This problem is based on the concept of chemical bonding.

VBT and VSEPR deal only with sigma bonds and valence electrons. There is no requirement of pi-bond in the process of determining the molecular shape.

Fundamentals

Theories such as VBT (valence bond theory) and VSEPR (valence shell electron pair repulsion theory) help in determining the shape, geometry and hybridization of a particular compound on the basis of various factors like valence bond, valence electrons and lone pairs.

Methane:

In case of methane (CH4)\left( {{\rm{C}}{{\rm{H}}_{\rm{4}}}} \right) molecule, the steric number is four in which the number of sigma bonds is 4 and number of lone pair of electrons at central atom is zero. Thus, bent, trigonal pyramidal, trigonal planar and linear molecular shapes are not possible.

The molecular shape of the methane is given below:

-С—н
Н-
Н
Н
Н
tetrahedral
І—0:
І—0—I

Carbonate ion:

In case of carbonate (CO32)\left( {{\rm{C}}{{\rm{O}}_3}^{ - 2}} \right) molecule, the steric number is 3 in which the number of sigma bonds is 3 and number of lone pair of electrons at central atom is zero. Thus, bent, trigonal pyramidal, linear and tetrahedral molecular shapes are not possible.

The molecular shape of carbonate is given below:

trigonal planar

Carbon dioxide:

In case of carbon di oxide (CO2)\left( {{\rm{C}}{{\rm{O}}_2}} \right) molecule, the steric number is 2 in which the number of sigma bonds is 2 and number of lone pair of electrons at central atom is zero. Thus, bent, trigonal pyramidal, trigonal planar and tetrahedral molecular shapes are not possible.

The molecular shape of carbon di oxide is given below:

о—с—
C=
Linear

Sulfite ion:

In case of sulfite (SO32)\left( {{\rm{S}}{{\rm{O}}_3}^{ - 2}} \right) molecule, the steric number is 3 in which the number of sigma bonds is 3 and number of lone pair of electrons at central atom is zero. Thus, bent, trigonal planar, linear and tetrahedral molecular shapes are not possible.

The molecular shape of sulfite is given below:

S
trigonal pyramidal

Ans:

The molecular shape of the methane molecule is tetrahedral.